Explain why the correct solution to the equation square root of x + 2 = x - 4 (which can be typed as sqrt(x + 2) = x – 4 ) is just 7. That is, explain what happens as you solve this equation and why you end up with the only solution x = 7.
Answer by richwmiller(17219)

sqrt(x + 2) = x – 4 )
square both sides
x+2=(x-4)^2
x+2=(x-4)^2
x+2=(x-4)*(x-4)
x+2=x^2-8x+16
0=x^2-9x+14
(x-7)* (x-2)=0
x=7 x=2
now why does 2 work
plug it back in the original.
sqrt(4)=0
2=0
But 2 doesn't equal 0

sqrt(7+2)=7-4
3=3